Please help me to find the inverse function of f in
problem f(x)=x^(2 ) ,x≥0
Thanks

Answer by nabla(475)

rewrite as:
y=x^2

Now swap x and y
x=y^2
Now solve for y
+/- Sqrt(x)=y
So f-1(x)=+/- Sqrt(x)